40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10933     Entry time: Fri Jan 23 02:11:40 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: CARM filters modified slightly 

[Jenne, Diego]

One of tonight's goals was to tweak the CARM filters, so that we could engage the lowpass filter, to avoid the detuned double cavity pole resonance disturbing the CARM loop.

I increased the Q of the zeros in the FM3 boost so that it eats fewer than the original 18 degrees of phase at 100 Hz.  We kept losing lock though, so for each lock I backed off on the Q a little bit.  In the end, the filter eats 9 degrees of phase at 100 Hz.  I also moved the lowpass from 700 Hz to 1kHz, although that doesn't change the phase at 100 Hz very much.

We modified the carm_up script re: PRMI locking a little bit.  The PRMI is not so enthusiastic about locking immediately at 25% MICH fringe, so I backed that off.  It now acquires lock at a few percent, and then ramps up the offset.  Also, the MICH FM6 bounce roll filter is now turned on after lock is acquired, effectively giving it an extra second or two of delay beyond the rest of the filters.

We were able several times to get to some MICH offset and turn on the lowpass filter, but starting to reduce the CARM offset makes us lose lock.  I think the problem is that the UGF servo demod phase is changing as we are changing offsets, filters and error signals.  We see that the I-phase is servoed successfully to 0dB, but that the Q-phase is starting to move around by 30 degrees or more.  We either need to monitor this much more closely, and add the changing demod phases to the carm_up script, or we need to go back to the sum-of-squares situation that we had last week.  Note that we saw failures with that method for a completely separate reason:  we were getting too close to the limiters, which cause the UGF servos to glitch and the outputs jump by a significant amount.  So, the issues that we were seeing last week when we had the sum-of-squares were a different thing, that we noticed and understood later.

Anyhow, nothing too exciting and glorious tonight, but progress has been made.

Also, from some Mist simulations that Q did, Diego made a sweet plot that is now posted in the control room, so we can translate arm power to CARM offset, at various MICH offsets. 

We also took some CARM loop measurements with the new filters.  We have a little more phase than we used to, which is nice.  These traces don't have the lowpass engaged, since I was trying to see how far we could get without it.  We lost lock right after the second measurement, but I think that was to do with the UGF servos.

Attachment 2: CARM_22Jan2015.pdf  19 kB  | Hide | Hide all
CARM_22Jan2015.pdf
ELOG V3.1.3-